OpenStudy (anonymous):

A line passes through (2,1) with slope 3. What is the equation of the line in slope-intercept form?

OpenStudy (anonymous):

Hey there, how are you?

OpenStudy (anonymous):

A slope-intercept form is what you call y=mx+b

OpenStudy (anonymous):

\[y=mx+b\] from here, you plug in x and y values that has been given .

OpenStudy (anonymous):

because coordinates come (x,y) you would end up doing this.

OpenStudy (anonymous):

\[(1)=3(2)+b\]

OpenStudy (anonymous):

So what would the equation look like?

OpenStudy (anonymous):

the equation will always look like y=mx+b

OpenStudy (anonymous):

because y=mx+b IS the slope intercept form

OpenStudy (anonymous):

So its y=3(2+1)

OpenStudy (anonymous):

no no

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

im sorry my words are wrong

OpenStudy (anonymous):

you plug in the x and y value to find out the y intercept which is b

OpenStudy (anonymous):

A line passes through (2,1) with slope 3. What is the equation of the line in slope-intercept form? y = 1 x = 2 correct?

OpenStudy (anonymous):



OpenStudy (anonymous):